What is the value of Kp of this reaction at 69.25 °C?



An important industrial process for synthesizing the ammonia used in agricultural fertilizers involves the combination of N2 and H2 

N2(g) +3H2(g) ⇌ 2NH3(g)  Kc= 3.654 x 108 at 69.25 °C 


What is the value of Kp of this reaction at 69.25 °C? 
Answer #1

Consider a given reaction N 2 (g) + 3 H 2 (g) 2 NH 3 (g)  

For this reaction, n = number of moles of gaseous products - number of moles of gaseous reactants

n = number of moles of gaseous NH 3 - [ No. of moles of gaseous N 2 + number of moles of gaseous H 2 ]

n = 2 - ( 1+3) = 2 - 4 = - 2

We have relation, K p = K c ( RT) n

Given : K c = 3.654 10 8 , R = 0.082057 L atm mol -1 K -1 , T = 69.25 + 273 = 342.25 K

K p = 3.654 10 8 ( 0.082057 342.25 ) - 2

K p = 4.633 105